8) Determine the normal boiling point of a substance whose vapor pressure is 55.1 mm Hg at 35 C 8) and has a AHvap of 32.1 kJmol B) 368 K A) 255 K C) 412 K D) 390. K E) 466 K 9) Which of the tollowing compounds is most soluble in hexane (CH3CH2CH2CH2CH2CH2CH3)? 9 A) ethanol 1-propanol C) 1-pentanol D) 1-butandl E) methanol

media%2F488%2F488fff56-b56e-4611-94ef-b6

media%2Fef1%2Fef10de67-2885-4f95-a5bf-9d
Answer #1

Ans 8

Clausius-Clapeyron equation

ln (P1/P2) =( \Delta Hvap /R) (1/T2 - 1/T1)

Normal boiling point occurs at 1 atm (760 mmHg)

P1 = 760 mmHg

P2 = 55.1 mmHg

T1 =?

T2 = 35+273 = 308 K

\DeltaHvap = 32.1*1000 J/mol

Gas constant R = 8.314 J/mol·K

ln (760/55.1) =( 32.1*1000 /8.314) (1/308 - 1/T1)

2.624 = 3860.957 (0.0032467 - 1/T1)

0.000679624 = (0.0032467 - 1/T1)

0.002567 = 1/T1

T1 = 389.54 K = 390K

Option D is the correct answer